OpenStudy (anonymous):

if 52 =64 that 36 =?

OpenStudy (anonymous):

if 52 = 64 then i think 36 = 48 but it depends on what ?

OpenStudy (anonymous):

no its not correct answer

OpenStudy (anonymous):

then tell me the correct answer please ..

OpenStudy (anonymous):

option are ....24,47,44,28

OpenStudy (dumbcow):

52 = 64 ??? prove it :) i think its 24

OpenStudy (anonymous):

well....52 Mod 10= 5*10, 2*1 ;;52 mod 8= 6* 8, 4* 1

OpenStudy (anonymous):

similarly 36 mod 10 = 3*10, 6* 1;;36 mod 8 =4*8, 4*1

OpenStudy (anonymous):

thus answer is 44. i found currently

OpenStudy (dumbcow):

ahh so 64 is 52 in base 8 and 44 is 36 in base 8

OpenStudy (anonymous):

yeah